### v1

@@ Abstract
The rapid evolution of deep learning (DL) models and the ever-increasing size of available datasets have raised the interest of the research community in the always important field of vision-based hand gesture recognition (VHGR), and delivered a wide range of applications, such as sign language understanding and human-computer interaction using cameras. Despite the large volume of research works in the field, a structured and complete survey on VHGR is still missing, leaving researchers to navigate through hundreds of papers in order to find the right combination of data, model, and approach for each task. The current survey aims to fill this gap by presenting a comprehensive overview of this aspect of computer vision. With a systematic research methodology that identifies the state-of-the-art works and a structured presentation of the various methods, datasets, and evaluation metrics, this review aims to constitute a useful guideline for researchers, helping them to choose the right strategy for delving into a certain VHGR task. Starting with the methodology used for study selection, literature retrieval, and the analytical framing, the survey identifies and organizes key VHGR approaches using a taxonomy-based format in various dimensions such as input modality and application domain. The core of the survey provides an in-depth analysis of state-of-the-art techniques across three primary VHGR tasks: static gesture recognition, isolated dynamic gestures and continuous gesture recognition. For each task, the architectural trends and learning strategies are listed. Additionally, the study reviews commonly used datasets —emphasizing on annotation schemes— and evaluates standard performance metrics. It concludes by identifying major challenges in VHGR, including both general computer vision issues and domain-specific obstacles, and outlines promising directions for future research.
@@
Vision-Based Hand Gesture Recognition; Human-Robot Interaction; Human-Computer Interaction; Sign Language; Deep Learning; Gesture Recognition Taxonomy
@@ 1 Introduction
Understanding human behaviors—such as hand gestures [2], actions [156], and facial expressions [150]—is a key area of artificial intelligence research. There is growing interest in developing Intelligent User Interfaces (IUIs) that offer more natural and intuitive ways for humans to interact with machines, replacing traditional tools like the keyboard and mouse. This has led to increased focus on Hand Gesture Recognition (HGR) across various applications. For example, HGR is used in human-robot interaction in industrial settings (e.g., robotic arm control [134, 96]), healthcare (e.g., robot-assisted surgery [139]), and autonomous mobile robot navigation (e.g., UAVs [194] and UGVs [17, 105]). It also plays a key role in human-computer interaction, especially in AR/VR environments [32, 199, 5]. In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic sign language recognition systems, aiming to improve communication and inclusion for these communities [196, 26].
@@
Given the importance of the applications mentioned above, Hand Gesture Recognition (HGR) has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based. Vision-based HGR relies on visual data such as RGB and depth images, as well as features derived from them—like pose heatmaps or optical flow. Sensor-based HGR, on the other hand, uses data from devices such as wearable gloves or surface EMG (sEMG) sensors, and even non-wearable sources like acoustic signals. Researchers have also explored multi-modal methods that combine inputs from different sources. In this study, we focus on vision-based HGR, as sensor-based methods often involve high cost and complex setup, making them less practical for real-world use. Vision-based approaches better align with the goal of HGR: to enable more natural, unobtrusive, and accessible human-machine interaction.
Many of the published surveys consider both sensor-based and vision-based methods together [140, 162, 64, 6]. As a result, the analysis of the proposed methods in these surveys tends to be superficial, because there are so many methods to cover. Other review papers also present the DAQ devices used by researchers [43], further broadening the range of related publications. The undeniable importance of sign language recognition has led some researchers to conduct surveys dedicated exclusively to this topic [2, 155, 170, 145, 174]. In addition, many authors consider both deep learning (DL) and traditional machine learning (ML) methods, or in some cases only traditional ML approaches [174, 140, 43, 132, 64, 6], while others include outdated methods [162, 123, 132, 7, 3, 84]. Although this broader approach helps lay the foundational understanding of VHGR and facilitates a historical perspective, it often limits the focus on current state-of-the-art methods, which are arguably more critical for driving the field forward.
Some surveys aim to provide a quantitative overview of the field [132, 64, 3, 84], offering bibliometric analyses on various aspects of VHGR. While such surveys present a statistical and objective perspective and help identify major subfields, they fall short in advancing the field due to their limited examination of the existing methodologies.
In contrast, to the aforementioned surveys, this survey emphasizes very recent publications and provides a detailed analysis of their contributions, aiming to identify current challenges and innovations that could serve as a valuable starting point for future research. A comparison with existing reviews is presented in Table 1.
Table 1: Comparison of this survey with existing ones. ML refers to traditional Machine Learning/Computer Vision techniques.
@@
Publication Year Range # Papers Input Data Models Scope
[2] 2021 - - Vision-based DL CSLR
[123] 2021 2014-2020 98 Vision-based DL & ML General
[7] 2021 2014-2021 138 Sensor- & Vision-based DL SLR
[3] 2021 2001-2021 649 Vision-based DL & ML SLR
[43] 2022 - - Sensor- & Vision-based DL & ML SLR
[132] 2022 2000-2020 571 Sensor- & Vision-based DL & ML General
[84] 2022 2010-2021 80 Vision-based DL & ML SLR
[155] 2023 2015-2023 - Vision-based DL ISLR
[162] 2024 2014-2024 - Sensor- & Vision-based DL & ML General
[64] 2024 2018-2023 256 Sensor- & Vision-based ML General
[6] 2024 2015-2024 58 Sensor- & Vision-based DL & ML SLR
[174] 2024 - - Sensor- & Vision-based DL & ML SLR
[140] 2024 - - Sensor- & Vision-based DL & ML General
This survey 2025 2021-2024 290 Vision-based DL General
@@
More specifically, the survey focuses on vision-based methods for three key Hand Gesture Recognition (HGR) tasks, which are discussed in detail in the following sections: static HGR [208, 131], isolated dynamic HGR (IDHGR) [82, 68, 214, 60, 206], and continuous dynamic HGR (CHGR) [122, 27, 74]. Static HGR involves recognizing gestures from single frames or spatial data only. Isolated dynamic HGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns. Continuous dynamic HGR goes a step further by detecting multiple gestures within long, unsegmented video sequences. Each task presents its own unique challenges. For example, CHGR often relies on weak supervision, making optimization more difficult. At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
@@
Vision-based Hand Gesture Recognition (HGR) has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines [124]. Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification. Pre-processing often involved segmenting the hand region using skin color or other visual cues. Feature extraction relied on descriptors like Histogram of Oriented Gradients (HOG), while classification was usually performed using models such as Support Vector Machines (SVMs) or Hidden Markov Models (HMMs), depending on the task. With the rise of deep learning, these hand-crafted features were gradually replaced by learned representations, which offer better generalization. Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks. These models first capture short-term motion patterns and then aggregate them to understand long-term context [122, 73, 62, 75, 74]. Due to their superior performance, this study focuses primarily on deep learning-based methods.
Given the broad scope of vision-based Hand Gesture Recognition (HGR) within computer vision, it is crucial to identify the current state-of-the-art methods for the three tasks mentioned above, as well as the key challenges that have shaped the field’s development. This survey aims to 1) provide an overview of the field, enabling readers to quickly find the most relevant publications for their needs, 2) offer a detailed analysis of each method, highlighting their key contributions, and 3) present the state-of-the-art solutions, based on the previous analysis, and explain the main challenges that continue to drive research in HGR.
@@
    A structured and comprehensive methodology for reviewing the literature;
@@
The structure of this paper is organized to guide the reader through a comprehensive review of vision-based hand gesture recognition (VHGR). Section 2 presents the literature review methodology, outlining the systematic approach used to collect and evaluate relevant works. Section 3 introduces a taxonomy of VHGR methods based on diverse criteria, offering a structured classification of existing approaches. Sections 4, 5 and 6 delve into the three main VHGR tasks –—static, isolated dynamic, and continuous gesture recognition—– discussing representative methods and their underlying principles. Special attention is given to sign language recognition and semaphoric gestures in dedicated subsections. Section 7 presents the main datasets used for the evaluation of the presented methods and Section 8 summarizes the evaluation metrics commonly used in the field. Finally, Section 9 outlines key challenges in Human Gesture Recognition and highlights current solutions, while Section 10 concludes the survey by reflecting on recent trends and highlighting open challenges that shape the future of VHGR research.
@@ 2 Literature Review Methodology
The review begins with a clear definition of the research methodology, establishing a structured foundation for the comprehensive exploration of the VHGR landscape. This systematic approach is intended to serve as a valuable reference for researchers and practitioners seeking to navigate and contribute to the field. The initial step involved defining the scope of the study, which informed the formulation of inclusion and exclusion criteria used to guide the selection of relevant publications. Subsequently, targeted queries were conducted across major scientific databases, including Scopus and IEEE Xplore, to retrieve works aligned with the established criteria. The final stage entailed a thorough examination of the selected literature, with the objective of distilling the key innovations, contributions, and methodological insights of each publication. These elements are further elaborated in the following sections.
@@ 2.1 Survey Scope
  • Established evaluation metrics used to measure system effectiveness;
@@
  • Innovative solutions and future directions aimed at overcoming existing limitations.
@@ 2.2 Selection Criteria
  • Up-to-date: To reflect the current state of the field, priority was given to publications from 2021 onward, particularly those introducing new methods. Exceptions were made for seminal works—defined here as milestone contributions post-2015 with high citation counts—that have significantly shaped VHGR research. In the case of datasets, no strict publication date was imposed, but only those actively used by the community or released within the past four years were included. All survey papers considered were published after 2021.
  • Impact: To further focus on seminal and influential works, publications were sourced from highly cited venues (i.e. having an h5-index of 90 or above on Google Scholar), with a single exception for the IEEE International Conference on Automatic Face & Gesture Recognition, which is closely aligned with VHGR. In addition to the above, and aiming to provide the theoretical context of the study, a small number of milestone papers has also been referenced, though not analyzed in the main methods section.
@@ 2.3 Literature Search
To identify publications that satisfy the previously defined criteria, a comprehensive literature search was conducted using specialized queries across several well-established scientific databases, including Scopus, IEEE Xplore, ACM Digital Library, and arXiv for pre-prints. Google Scholar was also used to double-check for publications from major venues that might have been overlooked. The search queries were designed based on the previously outlined inclusion/exclusion criteria, utilizing specific commands (e.g., EXACTSRCTITLE, LIMIT-TO) and boolean operators (e.g., AND, OR, NOT) to refine the results. A list of the keywords that were applied in the title, abstract and index terms of each one publication, can be found in Table 2. This process retrieved over 900 publications in the Scopus database. Out of them, only 210 papers were related to vision-based HGR. A second query has been performed on IEEE Xplore, excluding the non-IEEE venues, and resulted in 633 articles, most of them already listed in Scopus. A few simpler queries were performed on Google Scholar and arXiv, and duplicates were ignored, resulting in 290 papers in total, including review papers, methods, and datasets.
Table 2: Keywords used in literature search
Keywords
gesture recognition, gesture classification, sign language recognition, sign language understanding, pointing recognition, human finger pointing, hand gestures, human gestures, gesture, hand gesture, body gesture

### v2

@@ Abstract
The rapid evolution of deep learning (DL) models and the ever-increasing size of available datasets have raised the interest of the research community in the always important field of visual hand gesture recognition (VHGR), and delivered a wide range of applications, such as sign language understanding and human-computer interaction using cameras. Despite the large volume of research works in the field, a structured and complete survey on VHGR is still missing, leaving researchers to navigate through hundreds of papers in order to find the right combination of data, model, and approach for each task. The current survey aims to fill this gap by presenting a comprehensive overview of this computer vision field. With a systematic research methodology that identifies the state-of-the-art works and a structured presentation of the various methods, datasets, and evaluation metrics, this review aims to constitute a useful guideline for researchers, helping them to choose the right strategy for handling a VHGR task. Starting with the methodology used to locate the related literature, the survey identifies and organizes the key VHGR approaches in a taxonomy-based format, and presents the various dimensions that affect the final method choice, such as input modality, task type, and application domain. The state-of-the-art techniques are grouped across three primary VHGR tasks: static gesture recognition, isolated dynamic gestures, and continuous gesture recognition. For each task, the architectural trends and learning strategies are listed. To support the experimental evaluation of future methods in the field, the study reviews commonly used datasets and presents the standard performance metrics. Our survey concludes by identifying the major challenges in VHGR, including both general computer vision issues and domain-specific obstacles, and outlines promising directions for future research.
@@
Vision-Based Hand Gesture Recognition , Human-Robot Interaction , Human-Computer Interaction , Sign Language , Deep Learning , Gesture Recognition Taxonomy
[…]
@@ 1 Introduction
The recognition and interpretation of hand gestures (ADADOGLOU) play a central role in advancing Human–Computer Interaction (HCI) and assistive communication systems. In recent years, deep learning–based visual hand gesture recognition (VHGR) has gained significant momentum as an active subfield of computer vision, building upon progress in related areas such as action (survey-HAR-introduction-shafizadegan2024multimodal) and facial expression recognition (survey-FER-JESWA-2024). This has led to increased focus on Hand Gesture Recognition (HGR) across various applications. For example, HGR is widely used in Human-Robot Interaction (HRI) in industrial settings (e.g. robotic arm control (Papadopoulos2021Access; ROBOTIC_ARM)), in healthcare (e.g. robot-assisted surgery (TELEOPERATED_SURGICAL_ROBOT)), and in the autonomous navigation of mobile robots such as Unmanned Aerial (MD-UHGRD_SA-YOLO) and Ground (ULTRA-RANGE-UGV-Bamani; Linardakis2024Access) Vehicles (UAVs and UGVs). It also plays a key role in HCI, especially in AR/VR environments (OO-dMVMT; EgoGesture_DATASET; BOAT-MI_VR). In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic Sign Language Recognition (SLR) systems, aiming to improve communication and inclusion for these communities (survey-deaf-mute-AIR-2024; NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ).
@@
Given the importance of the applications mentioned above, HGR has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based. Vision-based HGR relies on visual data such as RGB and depth images, as well as features derived from them—like pose heatmaps or optical flow. Sensor-based HGR, on the other hand, uses data from devices such as wearable gloves or surface electromyography (sEMG) sensors, and even non-wearable sources like acoustic signals. Researchers have also explored multi-modal methods that combine inputs from different sources. In this study, we focus on vision-based HGR, as sensor-based methods often involve high cost and complex setup, making them less practical for real-world use. Vision-based approaches better align with the goal of HGR: to enable more natural, unobtrusive, and accessible gesture-based interaction.
Table 1: Comparison of this survey with indicative others. ML refers to traditional Machine Learning/Computer Vision techniques. The current survey mainly examines publications from the last four years, focusing on recent advancements regarding architectures, learning strategies, and related challenges, although older datasets that are relevant to this day have also been included.
@@
Publication Range Papers Input data Models Scope Comments Aspects that this survey aims to complement
(ADADOGLOU) 2016-2020 - Vision-based DL CSLR Provides an in-depth comparison of existing approaches, loss functions and pretraining schemes for CSLR. This survey expands the scope including non-CSLR methods and examines more recent methods.
(survey-5-9622242) 2014-2020 98 Vision-based DL & ML General Investigates the progress on hand gesture representation and Data Acquisition (DAQ) techniques as well as their performance and future research directions. It examines relatively antiquated methods. The current work aims to emphasize on recent DL approaches and related challenges and future perspectives.
(survey-10-al2021deep) 2014-2021 138 Sensor- & Vision-based DL SLR Categorizes Continuous and Isolated SLR methods based on their architectures and discusses pose estimation for ISLR in detail. This survey covers more topics including HCI & HRI, and highlights task-specific architectural patterns and challenges for CSLR.
(survey-11-adeyanju2021machine) 2001-2021 649 Vision-based DL & ML SLR Discusses relationships between research teams and DAQ techniques. It includes relatively old methods. This survey emphasizes on recently proposed ones in order to facilitate the evolution of the field.
(survey-6-ojeda2022survey) 2000-2020 571 Sensor- & Vision-based DL & ML General Provides many statistical results and bibliometric charts. It does not identify contemporary challenges or future research directions; this work seeks to detect relevant material.
(survey-3-el2022comprehensive) 1998-2020 - Sensor- & Vision-based DL & ML SLR Discusses in detail existing methods for SLR and provides multiple tables for quantitative comparison. This survey analyzes newer approaches for SLR and expands taxonomy and comparison to more application domains, beyond sign language.
(survey-16-joksimoski2022technological) 2010-2021 80 Vision-based DL & ML SLR A Natural Language Processing (NLP) toolkit was utilized to filter publications. Insightful statistics are also provided. Only a few methods are discussed and the current trends are not highlighted. On the contrary, this survey aims to identify trends and challenges.
(survey-17-sarhan2023unraveling) 2015-2023 - Vision-based DL ISLR Provides statistical charts and summary tables based on a four-fold taxonomy. This review introduces additional taxonomy criteria and covers more tasks.
(survey-4-shin2024methodological) 2014-2024 - Sensor- & Vision-based DL & ML General Considers a variety of modalities and examines methods for each one. The current work aims to discuss further multi-modality fusion strategies and identify more challenges.
(survey-8-hashi2024systematic) 2018-2023 256 Sensor- & Vision-based ML General Covers various aspects of HGR, including DAQ techniques, hand segmentation and feature extraction. Focuses exclusively on traditional ML approaches, which are currently surpassed by DL ones. This survey, searching for recent advancements, focuses on DL approaches.
(survey-9-al2024advancements) 2015-2024 58 Sensor- & Vision-based DL & ML SLR Examines various SLR methods and highlights the superiority of DL methods over older ones. Its does not identify challenges and future perspectives; this survey aims to cover this, going deeper into details.
(survey-13-tao2024sign) 2017-2024 - Sensor- & Vision-based DL & ML SLR Groups methods into RNN-based, CNN-based and transformers whilst discusses well-known benchmarks. This survey expands the scope to non-SLU papers and seeks to provide methods comparison and future prospects.
(survey-1-rahman2024comparative) 2008-2024 - Sensor- & Vision-based DL & ML General Covers a wide spectrum of gesture-based interaction and . The considered methods are relatively old. This survey aims to identify more recent advancements.
(linardakis2025_hgr_survey) 2018-2025 125 Vision-based DL & ML General Discuss jointly hand gesture capturing and recognition techniques. Many challenges are highlighted, including data bias and privacy issues. It covers a limited number of datasets for HGR and relatively old methods; this survey complements it by analyzing more datasets and recent advancements.
(survey-emporio2025continuous) 2002-2023 - Sensor- & Vision-based DL & ML CDHGR Focuses on online gesture recognition, discussing DAQ technologies and evaluation metrics, among others. This survey covers more topics and provides comparison of state-of-the-art methods.
(survey-khan2025deep) 2010-2024 - Sensor- & Vision-based DL & ML CSLR Discusses in detail methods and datasets for CSLR. It does not conduct methods comparison and the challenges are general; this survey highlights optimization difficulties in the context of CSLR.
The great importance and wide scope of HGR has led researchers to conduct surveys focusing on different aspects of the field, as shown in Table 1. Compared to existing review papers, this survey provides an in-depth examination of the recent Deep Learning (DL) methods (published since 2021), covering the current state-of-the-art (SoTA). In contrast, other existing reviews either focus on older approaches (published since 1998) (survey-6-ojeda2022survey; survey-3-el2022comprehensive; survey-8-hashi2024systematic) or mainly cover traditional machine learning techniques (survey-1-rahman2024comparative; survey-13-tao2024sign; linardakis2025_hgr_survey). While such works are valuable for understanding the fundamental methods and providing statistical insights, they are less suited to reflecting the recent evolution of the field. The present survey stands out by its broader and more up-to-date scope, including 238 papers published since 2021—-an average of nearly 47 publications per year. By comparison, (survey-11-adeyanju2021machine) and (survey-6-ojeda2022survey) analyze 649 and 571 papers respectively, but span roughly 20 years of research, corresponding to only about 30 publications per year. Finally, the current survey focuses exclusively on vision-based HGR. Including sensor-based approaches would significantly expand the scope and potentially confuse readers, while such methods generally offer limited expressiveness and reduced applicability in real-world scenarios. Nevertheless, the review maintains a broad and inclusive perspective, acknowledging that approaches from different domains may share common principles and can inform each other.
@@
More specifically, the survey focuses on vision-based methods for three key VHGR tasks, which are discussed in detail in the following sections: static HGR (SHGR) (FGDSNet; HAGRIDv2_ARXIV), isolated dynamic HGR (IDHGR) (SAM_SLR; SignBERT_ICCV; NLA_SLR; CTFB_Hampiholi; RAAR3DNet), and continuous (dynamic) HGR (CDHGR or CHGR) (VAC-CSLR; NEURIPS2022_6cd3ac24_HEATMAPS; CorrNet). SHGR involves recognizing gestures from single frames or spatial data only. IDHGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns. CDHGR goes a step further by detecting multiple gestures within long, unsegmented video sequences. Each task presents its own unique challenges. For example, CHGR often relies on weak supervision, making optimization more difficult. At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
@@
VHGR has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines (NVIDIA_GESTURE_R3DCNN_Molchanov). Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification. Pre-processing often involved segmenting the hand region using skin color or other visual cues. Feature extraction relied on descriptors like Histogram of Oriented Gradients (HOG), while classification was usually performed using models such as Support Vector Machines (SVMs) or Hidden Markov Models (HMMs), depending on the task. With the rise of deep learning, these hand-crafted features were gradually replaced by learned representations, which offer better generalization. Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks. These models first capture short-term motion patterns and then aggregate them to understand long-term context (VAC-CSLR; TemporalLiftPooling; SMKD; SEN-CSLR; CorrNet). Due to their superior performance, this study focuses primarily on deep learning-based methods.
@@
    A structured and comprehensive methodology for reviewing the VHGR literature;
@@
The structure of this paper aims to guide the reader through a comprehensive review of VHGR. Section 2 presents the literature review methodology, outlining the systematic approach used to collect and evaluate relevant works. Section 3 introduces a taxonomy of VHGR methods based on diverse criteria, offering a structured classification of existing approaches. Sections 4, 5 and 6 delve into the three main VHGR tasks –—static, isolated dynamic, and continuous gesture recognition—– discussing representative methods and their underlying principles. Special attention is given to sign language recognition and semaphoric gestures in dedicated subsections. Section 7 comprises a unified comparison of methods for VHGR regardless the task. Section 8 presents the main datasets used for the evaluation of the presented methods and Section 9 summarizes the evaluation metrics commonly used in the field. Finally, Section 10 outlines key challenges in Human Gesture Recognition and highlights current solutions, while Section 11 concludes the survey by reflecting on recent trends and highlighting open challenges that shape the future of VHGR research.
@@ 2 Literature Review Methodology
The literature review was conducted following a systematic methodology designed to ensure a comprehensive and structured exploration of the VHGR field. The study scope was first defined, guiding the development of inclusion and exclusion criteria for selecting relevant publications. Subsequently, targeted queries were executed across major scientific databases to retrieve works meeting these criteria. Finally, the selected literature was thoroughly examined to extract key innovations, contributions, and methodological insights, which are detailed in the following sections.
@@ 2.1 Survey Scope
  • 3.
    Established evaluation metrics used to measure the effectiveness of the proposed solutions;
  • 4.
@@
  • 5.
    Innovative solutions and future directions designed to address current limitations.
@@ 2.2 Selection Criteria
  • 2.
    Up-to-date: To capture the current state of the field, priority was given to publications from 2021 onward, especially those introducing new methods. Regarding datasets, no strict publication date was enforced; however, only those actively used by the community or released within the past four years were included. All surveyed review papers were published after 2021.
  • 3.
    Impact: To further emphasize seminal and influential works, publications were primarily sourced from highly cited venues (i.e. those with an h5-index of 90 or higher on Google Scholar), with a single exception for the IEEE International Conference on Automatic Face & Gesture Recognition, due to its strong relevance to VHGR. Additionally, to provide the theoretical context of the study, a small number of milestone papers were referenced, although they were not analyzed in detail in the main methods section.
@@ 2.3 Literature Search
To identify publications meeting the predefined criteria, a comprehensive literature search was performed using specialized queries across established scientific databases, including Scopus and Web of Science (WoS). Google Scholar was additionally consulted to capture potentially overlooked works from major venues, and arXiv was also used to find relevant high-impact pre-prints. Queries were constructed according to the inclusion/exclusion criteria, employing specific commands (e.g., EXACTSRCTITLE, LIMIT-TO) and Boolean operators (AND, OR, NOT) to refine results. The search targeted keywords in titles, abstracts, and index terms, including gesture recognition, gesture classification, sign language recognition, sign language understanding, pointing recognition, human finger pointing, hand gestures, human gestures, gesture, hand gesture, and body gesture. This process retrieved over 900 publications, of which 210 were relevant to vision-based HGR. Additional queries on Google Scholar and arXiv, with duplicates removed, produced a final set of 275 publications, encompassing review papers, methods, and datasets.